# DART: Difficulty-Adaptive Reasoning Truncation for Efficient Large Language Models

### v1

@@ 1 Introduction
The emergence of chain-of-thought (CoT) reasoning has marked a significant advance in enhancing the problem-solving abilities of LLMs by decomposing complex questions into intermediate steps (Wei et al., 2022; Kojima et al., 2022). Despite its effectiveness, the conventional CoT paradigm exhibits a critical inefficiency: it typically generates reasoning chains of a fixed, often excessive, length regardless of the inherent difficulty of the problem at hand (Chen et al., 2024; Fan et al., 2025; Sui et al., 2025). This “one-size-fits-all” approach results in substantial computational redundancy, increasing inference latency and resource consumption—a major bottleneck for deploying LLMs in applications.
@@
Adaptive reasoning, which aligns computational effort with problem difficulty, offers a promising path toward efficiency. Recent attempts to achieve such adaptability have largely relied on reinforcement learning (RL) frameworks, training models to penalize unnecessary reasoning length while preserving accuracy (Arora & Zanette, 2025; Ling et al., 2025; Zhang et al., 2025b; Shen et al., 2025). However, these RL-based methods remain unstable and heavily reward-dependent, suffering from high training difficulty and limited generalizability. Alternative approaches such as knowledge distillation focus on generating shorter reasoning chains (Yu et al., 2024; Kang et al., 2025; Xia et al., 2025), yet they produce statically compressed CoTs that lack dynamic adaptability. The key challenge is therefore to enable difficulty-aware reasoning in a stable and general manner, without relying on complex RL pipelines.
@@ 2 Related Work
Reasoning Chain Compression and Distillation. To address the computational overhead of lengthy reasoning chains, several compression and distillation techniques have been proposed. Knowledge distillation methods train smaller student models to mimic the reasoning processes of larger teacher models (Yu et al., 2024; Kang et al., 2025; Xia et al., 2025). These approaches typically produce statistically compressed reasoning chains that maintain a fixed length across all problems. Prompt-based compression techniques (Xu et al., 2025; Han et al., 2024) attempt to generate shorter reasoning chains through specialized prompting strategies, but often struggle with accuracy preservation on complex problems. Our work differs by introducing dynamic compression that adapts reasoning length to problem difficulty, achieving better efficiency-accuracy trade-offs than static compression methods.
Adaptive Reasoning Methods. Adaptive reasoning aims to dynamically adjust the reasoning process based on problem characteristics. Reinforcement learning approaches (Arora & Zanette, 2025; Yu et al., 2025; Shen et al., 2025; Zhang et al., 2025a) train length policies to decide when to halt reasoning, but face challenges with training stability and reward engineering. Prompt-based adaptive methods (Han et al., 2024) use heuristic rules to control reasoning length, but lack learning-based optimization. Unlike these approaches, DART employs a novel supervised learning framework that learns optimal reasoning length policies from automatically curated data, avoiding the instability of RL methods while maintaining architectural flexibility.
@@ 3.1 Motivation
To quantify this, we conducted a preliminary experiment on a subset of 10 problems from the MATH-500 (Lightman et al., 2023) dataset. Using our model fusion technique (detailed in Section 3.3), we generated reasoning chains for these problems across a spectrum of lengths, controlled by the fusion coefficient $\alpha$. For each problem, we evaluated the correctness of the answer yielded by chains of varying lengths. We then analyzed the aggregate accuracy as a function of the average number of tokens in the reasoning chains.
@@ 3.2 Step 1: Distilling Short CoTs
Base Model ($M_{\mathrm{base}}$). We select an efficient, open-source base model that natively supports chain-of-thought reasoning, such as DeepSeek-R1 (Guo et al., 2025) series or Qwen3 (Yang et al., 2025) series. This model, denoted as $M_{\mathrm{base}}$, is employed as-is for its ability to generate detailed, step-by-step reasoning. It serves as the efficient backbone whose behavior we aim to adapt, representing the long-chain reasoning style.
Short CoT Data Distillation ($\mathcal{D}_{\mathrm{short}}$).
We construct the short-chain dataset by compressing existing high-quality CoT data. Using a powerful distillation teacher model (e.g., DeepSeek-R1), we shorten each long reasoning chain $\mathrm{CoT}_{i}^{\mathrm{long}}$ to a concise version $\mathrm{CoT}_{i}^{\mathrm{short}}$ while preserving logical correctness:
@@ 3.4 Step 3: Curating the Adaptive Training Data
3. Optimal Chain Selection: From all models that produced the correct answer ($y_{i}^{\alpha}=y_{i}$), we select the reasoning chain $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ from the model with the largest value of $\alpha$ (i.e., the model biased towards the shortest chains). Formally: $\alpha^{*}=\min\{\alpha\in S\mid y_{i}^{\alpha}=y_{i}\}$, where $S$ is the set of sampled $\alpha$ values. The corresponding $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ is deemed the optimal adaptive chain for problem $x_{i}$.
@@ 4.1 Experimental Setup
Datasets. We evaluate our method on five challenging mathematical reasoning benchmarks: GSM8K (Cobbe et al., 2021), MATH-500 (Lightman et al., 2023), AMC23 (Committees, 2023), OLYMPAID (He et al., 2024), and AIME25 (Committees, 2025). These datasets span a wide range of difficulty levels, providing comprehensive evaluation of reasoning capabilities.
@@
Baselines. We also present a comprehensive comparison between our proposed method and several state-of-the-art reasoning optimization approaches. We categorize the baseline approaches into the following groups: (1) Prompt-based methods: These rely on prompt engineering without parameter updates. We select CoD (Xu et al., 2025) and TALE-EP (Han et al., 2024) as our baselines. (2) RL-based methods: These leverage reinforcement learning, typically with reward signals derived from task-specific objectives or human feedback, to optimize reasoning performance. We select Z1 (Yu et al., 2025), DAST (Shen et al., 2025), AdaptThink (Zhang et al., 2025a) as our baselines. (3) SFT-based methods: These utilize supervised fine-tuning on curated datasets to learn reasoning patterns. We select AutoL2S (Luo et al., 2025) as our baseline.
Implementation Details. We obtained Long CoT data from DeepSeek-R1-Distill(tuanha1305, 2025) dataset, which contains problem-solution pairs with detailed reasoning chains, and the short CoT compressed from the long chains using DeepSeek-R1 model with specific compression prompts. The complete prompt is provided in Appendix LABEL:A:prompt. We use Qwen3-14B as our base model to train a distilled version for model fusion. The training details can be refered in Appendix LABEL:A:train_full and the model fusion details can be refered in Appendix A.1. Adaptive training data is constructed from the training splits of GSM8K (Cobbe et al., 2021) and MATH datasets using our proposed curation pipeline. Please refer to Appendix A.2 for adaptive training details.
@@ 4.2 Main Results
Consistent Efficiency Gains.
DART consistently reduces computational cost across all model scales and datasets. The efficiency gains vary with problem difficulty: on simpler datasets like GSM8K, token usage is reduced by 34.0%–81.2%; while on highly challenging benchmarks AIME25, the method still achieves substantial savings up to 34.2%, demonstrating its ability to adaptively allocate more tokens to harder problems. This validates our core hypothesis that optimal reasoning length should adapt to problem difficulty.
@@ 4.3 Comparison to Prior Work
As shown in Table 1, our experimental analysis reveals a significant methodological divide in existing adaptive reasoning approaches. reinforcement learning (RL)-based methods (Z1, DAST, AutoL2S) demonstrate strong model specialization but are predominantly developed and evaluated exclusively on DeepSeek-R1-Distill-Qwen-7B, with no available implementations for Qwen3 series models. Conversely, prompt-based approaches (CoD, TALE-EP) offer broader model compatibility but face fundamental limitations in accuracy-efficiency trade-offs. This landscape highlights the unique positioning of our method in overcoming these constraints.
Superior Model Compatibility and Flexibility. Unlike RL-based methods that struggle with Qwen3’s extensive RLHF training, DART demonstrates remarkable architectural agnosticism. It achieves consistent improvements across all tested models (Qwen3-4B/8B/14B and DeepSeek-R1), proving its adaptability to diverse model architectures. While prompt-based methods like CoD and TALE-EP show wider model adaptability than RL-based techniques, they incur substantial accuracy costs across all model sizes. For instance, CoD exhibits accuracy degradation up to 6.6% on DeepSeek-R1, while TALE-EP shows similar patterns on Qwen3 series.
@@ 4.4.1 Validation of Model Fusion Effectiveness
As shown in Fig. 3, we observe that the average length of the generated CoT is a smooth, monotonically decreasing function of $\alpha$. This confirms that our fusion method successfully creates a controllable knob for adjusting reasoning complexity without additional training. Furthermore, the accuracy on reasoning benchmarks reveals a more nuanced relationship: as $\alpha$ decreases from 1 to 0, accuracy initially shows a slight improvement before entering a declining phase. This non-monotonic behavior demonstrates that longer reasoning chains do not invariably lead to higher accuracy; beyond a certain point, excessive verbosity can be counterproductive. The fusion spectrum thus effectively samples different points on the efficiency-accuracy Pareto frontier, providing an efficient and elegant solution for generating the multi-length reasoning data required for the next step of our pipeline.
@@ 5 Conclusion
Appendix A Appendix
A.1 Details of Model Fusion
@@
A.2 Details of Adaptive Model Training
We perform adaptive fine-tuning on the model using the LLaMA-Factory (Zheng et al., 2024) framework with Low-Rank Adaptation (LoRA). The LoRA configuration targets all linear layers (lora_target: all) with a rank of 256 and an alpha of 16. The training source question data comprises about 15,000 samples from the GSM8K and MATH mixture dataset, using the qwen3 prompt template. A cutoff length of 32,768 tokens is applied to all sequences.

### v2

@@ 1 Introduction
The emergence of chain-of-thought (CoT) reasoning has marked a significant advance in enhancing the problem-solving abilities of LLMs by decomposing complex questions into intermediate steps (wei2022chain; kojima2022large). Despite its effectiveness, the conventional CoT paradigm exhibits a critical inefficiency: it typically generates reasoning chains of a fixed, often excessive, length regardless of the inherent difficulty of the problem at hand (chen2024not; fan2025missing; sui2025stop). This “one-size-fits-all” approach results in substantial computational redundancy, increasing inference latency and resource consumption—a major bottleneck for deploying LLMs in applications.
@@
Adaptive reasoning, which aligns computational effort with problem difficulty, offers a promising path toward efficiency. Recent attempts to achieve such adaptability have largely relied on reinforcement learning (RL) frameworks, training models to penalize unnecessary reasoning length while preserving accuracy (arora2025training; ling2025fast; zhang2025continue; shen2025dast). However, these RL-based methods remain unstable and heavily reward-dependent, suffering from high training difficulty and limited generalizability. Alternative approaches such as knowledge distillation focus on generating shorter reasoning chains (yu2024distilling; kang2025c3ot; xia2025tokenskip), yet they produce statically compressed CoTs that lack dynamic adaptability. The key challenge is therefore to enable difficulty-aware reasoning in a stable and general manner, without relying on complex RL pipelines.
@@ 2 Related Work
Reasoning Chain Compression and Distillation. To address the computational overhead of lengthy reasoning chains, several compression and distillation techniques have been proposed. Knowledge distillation methods train smaller student models to mimic the reasoning processes of larger teacher models (yu2024distilling; kang2025c3ot; xia2025tokenskip). These approaches typically produce statistically compressed reasoning chains that maintain a fixed length across all problems. Prompt-based compression techniques (xu2025chain; han2024token) attempt to generate shorter reasoning chains through specialized prompting strategies, but often struggle with accuracy preservation on complex problems. Our work differs by introducing dynamic compression that adapts reasoning length to problem difficulty, achieving better efficiency-accuracy trade-offs than static compression methods.
Adaptive Reasoning Methods. Adaptive reasoning aims to dynamically adjust the reasoning process based on problem characteristics. Reinforcement learning approaches (arora2025training; yu2025z1; shen2025dast; zhang2025adaptthink) train length policies to decide when to halt reasoning, but face challenges with training stability and reward engineering. Prompt-based adaptive methods (han2024token) use heuristic rules to control reasoning length, but lack learning-based optimization. Unlike these approaches, DART employs a novel supervised learning framework that learns optimal reasoning length policies from automatically curated data, avoiding the instability of RL methods while maintaining architectural flexibility.
@@ 3.1 Motivation
To quantify this, we conducted a preliminary experiment on a subset of 10 problems from the MATH-500 (lightman2023let) dataset. Using our model fusion technique (detailed in Section 3.3), we generated reasoning chains for these problems across a spectrum of lengths, controlled by the fusion coefficient $\alpha$. For each problem, we evaluated the correctness of the answer yielded by chains of varying lengths. We then analyzed the aggregate accuracy as a function of the average number of tokens in the reasoning chains.
@@ 3.2 Step 1: Distilling Short CoTs
Base Model ($M_{\mathrm{base}}$). We select an efficient, open-source base model that natively supports chain-of-thought reasoning, such as DeepSeek-R1 (guo2025deepseek) series or Qwen3 (yang2025qwen3) series. This model, denoted as $M_{\mathrm{base}}$, is employed as-is for its ability to generate detailed, step-by-step reasoning. It serves as the efficient backbone whose behavior we aim to adapt, representing the long-chain reasoning style.
Short CoT Data Distillation ($\mathcal{D}_{\mathrm{short}}$). We construct the short-chain dataset by compressing existing high-quality CoT data. Using a powerful distillation teacher model (e.g., DeepSeek-V3 (guo2025deepseek)), we shorten each long reasoning chain $\mathrm{CoT}_{i}^{\mathrm{long}}$ to a concise version $\mathrm{CoT}_{i}^{\mathrm{short}}$ while preserving logical correctness:
@@ 3.4 Step 3: Curating the Adaptive Training Data
3. Optimal Chain Selection: From all models that produced the correct answer ($y_{i}^{\alpha}=y_{i}$), we select the reasoning chain $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ from the model with the largest value of $\alpha$ (i.e., the model biased towards the shortest chains). Formally: $\alpha^{*}=\max\{\alpha\in S\mid y_{i}^{\alpha}=y_{i}\}$, where $S$ is the set of sampled $\alpha$ values. The corresponding $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ is deemed the optimal adaptive chain for problem $x_{i}$.
@@ 4.1 Experimental Setup
Backbone Reasoning Models. We conduct experiments on four representative open-source large language models with strong reasoning capabilities: the Qwen3 (yang2025qwen3) series (4B-Thinking-2507  [^1] Throughout this paper, all Qwen3-4B results refer to the Qwen3-4B-Thinking-2507 variant., 8B, and 14B parameter versions) and DeepSeek-R1-Distill-Qwen-7B (guo2025deepseek). These models were selected for their native chain-of-thought reasoning capabilities and represent diverse architectural approaches to reasoning in language models.
Datasets. We evaluate our method on five challenging mathematical reasoning benchmarks: GSM8K (cobbe2021training), MATH-500 (lightman2023let), AMC23 (amc23), OLYMPAID (he2024olympiadbench), and AIME25 (aime25). These datasets span a wide range of difficulty levels, providing comprehensive evaluation of reasoning capabilities.
@@
Baselines. We also present a comprehensive comparison between our proposed method and several state-of-the-art reasoning optimization approaches. We categorize the baseline approaches into the following groups: (1) Prompt-based methods: These rely on prompt engineering without parameter updates. We select CoD (xu2025chain) and TALE-EP (han2024token) as our baselines. (2) RL-based methods: These leverage reinforcement learning, typically with reward signals derived from task-specific objectives or human feedback, to optimize reasoning performance. We select Z1 (yu2025z1), DAST (shen2025dast), AdaptThink (zhang2025adaptthink) as our baselines. (3) SFT-based methods: These utilize supervised fine-tuning on curated datasets to learn reasoning patterns. We select AutoL2S (luo2025autol2s) as our baseline.
Implementation Details. We obtained Long CoT data from DeepSeek-R1-Distill(tuanha1305) dataset, which contains problem-solution pairs with detailed reasoning chains, and the short CoT compressed from the long chains using DeepSeek-R1 (guo2025deepseek) model with specific compression prompts. The complete prompt is provided in Appendix A.1. We use Qwen3-14B as our base model to train a distilled version for model fusion. The training details can be refered in Appendix A.2 and the model fusion details can be refered in Appendix A.3. Adaptive training data is constructed from the training splits of GSM8K (cobbe2021training) and MATH (hendrycks2021measuring) datasets using our proposed curation pipeline. Please refer to Appendix A.4 for adaptive training details.
@@ 4.2 Main Results
Consistent Efficiency Gains. DART consistently reduces computational cost across all model scales and datasets. The efficiency gains vary with problem difficulty: on simpler datasets like GSM8K, token usage is reduced by 34.0%–81.2%; while on highly challenging benchmarks AIME25, the method still achieves substantial savings up to 34.2%, demonstrating its ability to adaptively allocate more tokens to harder problems. This validates our core hypothesis that optimal reasoning length should adapt to problem difficulty.
@@ 4.3 Comparison to Prior Work
As shown in Table 1, our experimental analysis reveals a significant methodological divide in existing adaptive reasoning approaches. RL-based methods (Z1, DAST, AdaptThink) demonstrate strong model specialization but are predominantly developed and evaluated exclusively on DeepSeek-R1-Distill-Qwen-7B, with no available implementations for Qwen3 series models. Conversely, prompt-based approaches (CoD, TALE-EP) offer broader model compatibility but face fundamental limitations in accuracy-efficiency trade-offs. This landscape highlights the unique positioning of our method in overcoming these constraints.
Superior Model Compatibility and Flexibility. Unlike RL-based methods that struggle with Qwen3’s extensive RLHF training, DART demonstrates remarkable architectural agnosticism. It achieves consistent improvements across all tested models (Qwen3-4B/8B/14B and DeepSeek-R1), proving its adaptability to diverse model architectures. While prompt-based methods like CoD and TALE-EP show wider model adaptability than RL-based techniques, they incur substantial accuracy costs across all model sizes. For instance, CoD exhibits accuracy degradation up to 6.6 on DeepSeek-R1-Distill-Qwen-7B, while TALE-EP shows similar patterns on all tested models.
@@
4.4 Generalization to Non-Mathematical Reasoning Tasks
To further validate the generalizability of the DART framework, we evaluate its performance on three challenging non-mathematical reasoning benchmarks: GPQA (rein2024gpqa), a demanding dataset of PHD-level questions in physics, biology, and chemistry; LogiQA (liu2020logiqa), a dataset for logical reasoning; and CommonsenseQA (talmor2019commonsenseqa), which tests commonsense reasoning. Crucially, the DART models evaluated here are the same models trained only on the mathematical reasoning datasets (GSM8K and MATH)–without any domain-specific fine-tuning–assessing their zero-shot cross-domain transfer capability.
As shown in Table 2, DART demonstrates robust cross-domain generalization with consistent efficiency improvements and accuracy preservation across all domains. On GPQA, DART reduces ACT by 16.6%-46.9% while improving Pass@1 accuracy by up to 5.1 (Qwen3-14B); on LogiQA, ACT reductions of 20.9%-48.5% are achieved while maintaining comparable accuracy; and on CommonsenseQA, efficiency gains of 7.0%-49.2% are accompanied by accuracy improvements of up to 10.6 (DeepSeek-R1-Distill-Qwen-7B). This consistent pattern across scientific, logical, and commonsense reasoning tasks demonstrates that DART learns fundamental principles of reasoning sufficiency rather than domain-specific patterns, effectively establishing its value as a general-purpose adaptive reasoning framework.
Table 2: Cross-domain generalization performance of DART on non-mathematical reasoning benchmarks. All DART models were trained only on mathematical reasoning datasets (GSM8K and MATH).
Method GPQA LogiQA CommonsenseQA
Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT
Qwen3-4B
Vanilla 63.1+0.0 8048.96-0.0% 9070.81-0.0% 81.6+0.0 3572.32-0.0% 4268.86-0.0% 71.6+0.0 2037.42-0.0% 2536.90-0.0%
DART(Ours) 66.2+3.1 6710.46-16.6% 7558.13-16.7% 81.4-0.2 2825.63-20.9% 3368.78-21.1% 77.0+5.4 1895.72-7.0% 2223.47-12.4%
Qwen3-8B
Vanilla 58.1+0.0 8632.48-0.0% 9660.64-0.0% 80.2+0.0 2533.89-0.0% 3127.05-0.0% 75.6+0.0 1141.44-0.0% 1621.51-0.0%
DART(Ours) 57.8-0.3 5980.81-30.7% 6790.24-29.7% 80.7+0.5 1689.21-33.3% 2222.22-28.9% 77.4+1.8 807.28-29.3% 1193.31-26.4%
Qwen3-14B
Vanilla 60.6+0.0 7161.52-0.0% 8083.38-0.0% 83.7+0.0 1970.96-0.0% 2544.78-0.0% 75.4+0.0 884.82-0.0% 1289.27-0.0%
DART(Ours) 65.7+5.1 5286.08-26.2% 6123.59-24.2% 84.5+0.8 1476.42-25.1% 2012.44-20.9% 77.4+2.0 663.28-25.0% 1028.92-20.2%
DeepSeek-R1-Distill-Qwen-7B
Vanilla 40.9+0.0 6166.27-0.0% 7158.12-0.0% 48.7+0.0 1633.06-0.0% 2196.57-0.0% 40.0+0.0 885.90-0.0% 1322.53-0.0%
DART(Ours) 43.9+3.0 3275.42-46.9% 5953.13-16.8% 48.7+0.0 840.55-48.5% 1474.76-32.9% 50.6+10.6 449.62-49.2% 953.00-27.9%
@@ 4.5.1 Validation of Model Fusion Effectiveness
As shown in Fig. 3, we observe that the average length of the generated CoT is a smooth, monotonically decreasing function of $\alpha$. This confirms that our fusion method successfully creates a controllable knob for adjusting reasoning complexity without additional training. Furthermore, the accuracy on reasoning benchmarks reveals a more nuanced relationship: as $\alpha$ increases from 0 to 1, accuracy initially shows a slight improvement before entering a declining phase. This non-monotonic behavior demonstrates that longer reasoning chains do not invariably lead to higher accuracy; beyond a certain point, excessive verbosity can be counterproductive. The fusion spectrum thus effectively samples different points on the efficiency-accuracy Pareto frontier, providing an efficient and elegant solution for generating the multi-length reasoning data required for the next step of our pipeline.
@@ 4.5.2 Impact of Data Curation Repetition
4.5.4 Explicit Validation of Difficulty-Aware Adaptation.
To explicitly validate DART’s ability to adapt reasoning length based on problem difficulty, we conduct a fine-grained analysis using the MATH-500 dataset with explicit difficulty ratings (level 1-5). As illustrated in Figure 5, evaluating DART (Qwen3-4B) across these levels reveals a clear difficulty-aware pattern: for easier problems (levels 1-2), high compression ratios (56.9% and 56.4%) are achieved while maintaining near-perfect accuracy (100.0% and 96.7%). As difficulty increases to level 3, compression slightly decreases to 55.2% to preserve accuracy (98.1%), and for the hardest problems (levels 4-5), compression further drops to 47.1% and 34.0%, prioritizing accuracy over efficiency.
Figure 5: Accuracy vs. compression ratio across different difficulty levels on MATH-500 dataset (Qwen3-4B).
This graded adaptation demonstrates that DART successfully internalizes the mapping from problem difficulty to optimal reasoning length. Rather than applying uniform compression, it dynamically allocates more computational resources to harder problems while truncating reasoning on simpler ones, confirming its ability to make nuanced decisions based on problem complexity.
@@ 5 Conclusion
Appendix A Implementation Details
A.1 Prompt Template of Short CoT Data Distillation
Figure 6: Prompt template of short CoT data distillation.
The full prompt for prompt template of short CoT data distillation is shown in Figure 6.
A.2 Details of Distilled Model Training
We fine-tune the Qwen3-14B model using the LLaMA-Factory (zheng2024llamafactory) framework, employing a full-parameter fine-tuning approach. The training process is optimized with the DeepSpeed ZeRO Stage 3 strategy. The training data consists of 30,000 math samples from the DeepSeek-R1-Distill(tuanha1305) dataset. We set a cutoff length of 32,768 tokens for the sequences.
The model is fine-tuned for 3 epochs with the following hyperparameters:
  • Cutoff length: 32,768
  • Max samples: 30,000
  • Batch size: 1 (with a gradient accumulation of 2)
  • Learning rate: $5\times 10^{-6}$ with a cosine schedule and a warmup ratio of 0.1
  • Precision: bf16
  • Validation split: 10% of the training data
  • Evaluation strategy: every 200 steps
All experiments are conducted with the overwrite_cache=true option and utilize 16 parallel workers for data preprocessing. The resulting models are directly used for downstream evaluation without any additional tuning.
A.3 Details of Model Fusion
@@
A.4 Details of Adaptive Model Training
We perform adaptive fine-tuning on the model using the LLaMA-Factory (zheng2024llamafactory) framework with Low-Rank Adaptation (LoRA) (hu2022lora). The LoRA configuration targets all linear layers (lora_target: all) with a rank of 256 and an alpha of 16. The training source question data comprises about 15,000 samples from the GSM8K and MATH mixture dataset, using the qwen3 prompt template. A cutoff length of 32,768 tokens is applied to all sequences.